# SMILES & SELFIES have to go: Representation of Molecules via Algebraic Data Types

### v2

@@ 5.4.1 Example of Orbitals : Hydrogen, Carbon & Oxygen
1hydrogen :: Shells
2hydrogen =
@@
5 , sSubShell = Just (SubShell [Orbital So 1 Nothing]) -- 1s^1
@@
11
12carbon :: Shells
13carbon =
@@
16 , sSubShell = Just (SubShell [Orbital So 2 Nothing]) -- 1s^2
@@
23 , sSubShell = Just (SubShell [Orbital So 2 Nothing]) -- 2s^2
24 , pSubShell = Just (SubShell [Orbital Px 1 (Just (Coordinate 1 0 0)),
25 Orbital Py 1 (Just (Coordinate 0 1 0))])
26 -- 2p^2
@@ 25 Orbital Py 1 (Just (Coordinate 0 1 0))])
31
32oxygen :: Shells
33oxygen =
34 [ Shell
35 { principalQuantumNumber = 1
36 , sSubShell = Just (SubShell [Orbital So 2 Nothing]) -- 1s^2
37 , pSubShell = Nothing
38 , dSubShell = Nothing
39 , fSubShell = Nothing
@@
41 , Shell
42 { principalQuantumNumber = 2
43 , sSubShell = Just (SubShell [Orbital So 2 Nothing]) -- 2s^2
44 , pSubShell = Just (SubShell [Orbital Px 2 (Just (Coordinate 1 0 0)),
45 Orbital Py 1 (Just (Coordinate 0 1 0)),
46 Orbital Pz 1 (Just (Coordinate 0 0 1))])
47 -- 2p^4
48 , dSubShell = Nothing
49 , fSubShell = Nothing
@@ 46 Orbital Pz 1 (Just (Coordinate 0 0 1))])
51 ]
Listing 8: Examples of electronic configurations for hydrogen, carbon, and oxygen using the Orbital Algebraic Data Type representation in Haskell.
For the hydrogen atom, the electronic configuration is represented by a single shell with a principal quantum number of 1. The shell contains only an $s$ subshell (sSubShell) with one orbital occupied by a single electron, denoted as Orbital So 1 Nothing. The Nothing value for the orientation indicates that s orbitals have a spherical symmetry and do not have a specific orientation. This represents the ground state configuration of hydrogen, which is 1s1.
The carbon atom’s electronic configuration is represented by two shells. The first shell, with a principal quantum number of 1, consists of an $s$ subshell with two electrons (Orbital So 2 Nothing), corresponding to the 1s2 configuration. The second shell, with a principal quantum number of $2$, comprises an $s$ subshell with two electrons (Orbital So 2 Nothing) and a $p$ subshell (pSubShell) with two orbitals, each containing one electron. The p orbitals are specified as Orbital Px 1 (Just (Coordinate 1 0 0)) and Orbital Py 1 (Just (Coordinate 0 1 0)), representing the 2px and 2py orbitals, respectively. The Coordinate values (1, 0, 0) and (0, 1, 0) indicate the orientation of the p orbitals along the x and y axes, respectively. This configuration, 1s22s22p2, describes the ground state of carbon.
Similarly, the oxygen atom’s electronic configuration is represented by two shells. The first shell is identical to that of carbon, with an $s$ subshell containing two electrons. The second shell consists of an $s$ subshell with two electrons and a $p$ subshell with four electrons distributed across three orbitals: Orbital Px 2 (Just (Coordinate 1 0 0)), Orbital Py 1 (Just (Coordinate 0 1 0)), and Orbital Pz 1 (Just (Coordinate 0 0 1)). The Coordinate values (1, 0, 0), (0, 1, 0), and (0, 0, 1) represent the orientation of the $p$ orbitals along the $x$, $y$, and $z$ axes, respectively. This describes the ground state configuration of oxygen, which is 1s22s22p4. The oxygen atom’s electronic configuration differs from carbon in the occupancy of the p subshell, where oxygen has four electrons distributed across three p orbitals, while carbon has two electrons in two p orbitals.
The inclusion of the Coordinate type in the representation allows for the specification of the spatial orientation of the orbitals. This is particularly relevant for $p$, $d$, and $f$ orbitals, which have distinct orientations in three-dimensional space. By providing the coordinates, the representation captures the directional nature of these orbitals, enabling a more accurate description of the electronic structure and the spatial distribution of electrons in the atom.
6 Experiments and extensions
6.1 Reaction Networks
The proposed algebraic data type for representing molecules can be extended to model chemical reactions and reaction networks. By defining a Reaction data type that captures the reactants, products, and conditions under which a reaction occurs, we can create a powerful tool for studying and simulating chemical processes. A chemical reaction network comprises a set of reactants, a set of products, and a set of reactions.
Reactions are commutative but not associative and given further time could be incorporated into the type. This would be called an abelian or commutative magma.
Here’s an example of how the Reaction data type can be defined: 1data Reaction = Reaction { 2 reactants :: [(Double, Molecule)] 3 , products :: [(Double, Molecule)] 4 , conditions :: [Condition] 5 , reactionRate :: Double 6} 7 8data Condition = TempCondition {temperature :: Double, times :: Times} 9 | PressureCondition {pressure :: Double, times :: Times} 10 11newtype Times = Times (StartTime Double, EndTime Double) Listing 9: Haskell representation of a chemical reaction including reactants, products, conditions, and reaction rate.
In this representation, a Reaction consists of a list of reactants and products, each represented as a tuple of a stoichiometric coefficient (of type Double) and an Atom. The conditions field is a list of Conditions, which can be either temperature or pressure conditions, specified along with the time range over which they apply.
Using this Reaction data type, we can define a ChemicalReactionNetwork as a collection of reactions:
1newtype ChemicalReactionNetwork = ChemicalReactionNetwork {
2 reactions :: [Reaction]
3}
Listing 10: Haskell representation of a chemical reaction including reactants, products, conditions, and reaction rate.
This allows for the modeling of complex chemical systems and the study of their behavior under different conditions. For example, consider the reaction for the formation of water from hydrogen and oxygen:
1exampleReaction :: Reaction
2exampleReaction = Reaction {
3 reactants = [(2.0, h2), (1.0, o2)],
4 products = [(2.0, h2o)],
5 conditions = [ TempCondition 500.0 (Times (0.0, 10.0)),
6 PressureCondition 1.0 (Times (0.0, 10.0)) ],
7 reactionRate = 0.1
8}
Listing 11: Haskell representation of a chemical reaction including reactants, products, conditions, and reaction rate.
This reaction represents the formation of two water molecules from two hydrogen molecules and one oxygen molecule, under the conditions of a temperature of $500K$ and a pressure of $1$ atm, over a time range of 00 to $10$ seconds.
@@ 6.2 Group Theory
Here’s an example of how a MoleculeRotation data type can be defined and made an instance of the Group typeclass: 1class Group g where 2 mul :: g -> g -> g -- Group operation: combining two operations 3 inv :: g -> g -- Inverse of an operation 4 e :: g -> g -- Identity operation 5 6data MolRot = MolRot Molecule Coordinate Double deriving (Show) 7 8instance Group MolRot where 9 mul (MolRot mol1 axis1 angle1) (MoleculeRotation mol2 axis2 angle2) = 10 combineRotations mol1 axis1 angle1 axis2 angle2 11 12 inv (MoleculeRotation mol axis angle) = MolRot mol axis (-angle) 13 14 e (MoleculeRotation mol axis angle) = MolRot mol (Coordinate 0 0 0) 0 Listing 12: Haskell representation of a chemical reaction including reactants, products, conditions, and reaction rate. In this representation, a MolRot consists of a Molecule, a rotation axis specified by a Coordinate, and a rotation angle (in radians). The Group typeclass defines the necessary operations for a group, namely the group operation (mul), the inverse operation (inv), and the identity operation (e).
@@ 6.3 Storage, Transmission, and Readability
As well the computational purposes discussed elsewhere in the paper, the ADT provides advantages with respect to storage of the molecule, and transmission between cheminformaticians.
Molecules in the ADT can be serialised and stored as .hs files, which may be used for storage, and for transmission of molecular data, perhaps via large chemical databases, through straightforward use of the Show and Read instances in Haskell, as shown in Listing 14.
Additionally, as .hs files will be interpreted with Haskell, molecules can be straightforwardly compressed as Listing 13 demonstrates by using foldr to compress the bondsData list of the benzene molecule. Although this may raise the issue of “canonicalisation” for storage of molecules, it should be noted the bondsData in Listing 13 and Listing 6 will be stored the same way inside of a Haskell program, as the corresponding data is the same in each case. This could be beneficial in the representation of large molecules, as only a small number of bonds or atoms need be explicitly described to encode the molecule for storage.
1
2bondsData :: [((Int, Int), Bond)]
@@
Listing 13: Using foldr to compress the bondsData of the molecule benzene in the ADT representation. As Benzene is symmetric, the bonding system can be specified by only describing three bonding systems, and using those to construct the bondsData list.
@@
Listing 14: Haskell code demonstrating the serialisation and writing to disk of a Molecule data type and the reading of a molecule stored in an .hs file, through use of the show and read functions.
An example of the usage is shown.
@@
1-- Observed molecule (e.g., H2O)
2observedMolecule :: Molecule
3observedMolecule = h2o
4
5-- A generative model for a molecule
6moleculeModel :: Meas Molecule
7moleculeModel = do
@@
9 atoms <- replicateM numAtoms \$ do
@@
13 let shellsVar = elementShells symbol
14 return \$ Atom { atomID = 0,
15 atomicAttr = attr,
16 coordinate = coord,
17 shells = shellsVar }
18
19 let molecule = Molecule { atoms = atoms, bonds = randomBonds 3 }
20
21 -- Score the model based on the distance to the observed molecule
22 let distance = distMolecule molecule observedMolecule
@@
26
27-- Sample a coordinate from a normal distribution
28sampleCoordinate :: Meas Coordinate
29sampleCoordinate = do
@@
34
35-- Calculate the Euclidean distance between two coordinates
36euclideanDistance :: Coordinate -> Coordinate -> Double
37euclideanDistance (Coordinate x1 y1 z1) (Coordinate x2 y2 z2) =
@@
39
40-- Calculate the distance between two molecules
41distMolecule :: Molecule -> Molecule -> Double
42distMolecule mol1 mol2 =
@@
45 distances = [euclideanDistance c1 c2 | c1 <- coords1, c2 <- coords2]
46 in minimum distances
47
48-- Example usage
49main :: IO ()
50main = do
51 samples <- mh 0.1 moleculeModel
52 let (molecules, weights) = unzip \$ take 1000 \$ drop 1000 samples
53 putStrLn \$ "Sampled-molecules:" ++ show molecules
54 putStrLn \$ "Weights:" ++ show weights
Listing 15: In a probabilistic programming context, trace Metropolis-Hastings (M-H) operates by sampling entire “traces” of latent variables (i.e., all random choices) from a model [42]. Each step proposes a small perturbation to the current trace, controlled by a jitter parameter, often the standard deviation of the proposal distribution. This parameter must be tuned carefully to ensure efficient exploration of the posterior while maintaining a reasonable acceptance rate. Typically, an initial burn-in period is applied, during which early samples (which have not yet converged) are discarded. This burn-in helps remove the influence of the initial point on the chain, leading to samples that more accurately reflect the stationary distribution of the target posterior. Here the jitter is set to 0.1 and the number of burnin samples is 1000.
6.4 A simple example of probablistic programming
The code block above can be explained as follows:
  • We define observedMolecule as the molecule we want to condition our model on. In this example, we use the predefined h2o molecule, but you can replace it with any other molecule of interest.
  • We define a generative model moleculeModel using the Meas monad. This model generates a Molecule by first sampling the number of atoms numAtoms from a discrete uniform distribution using uniformD. In this case, we set numAtoms to a fixed value of 3.
  • We then use replicateM to generate numAtoms number of Atoms. For each Atom, we sample the symbol uniformly from the list [C, N, O, H] using uniformD, and sample the coordinate using the sampleCoordinate function.
  • We use record syntax to create each Atom. The atomID is set to a default value of 0 (which can be updated later), the atomicAttr is set using the elementAttributes function based on the sampled symbol, and the shells are set using the elementShells function, also based on the symbol. The charge is assumed to be neutral at this point.
  • After generating the list of atoms, we create a Molecule using record syntax. The atoms field is set to the generated list of atoms, and the bonds field is set using the randomBonds function, which randomly initializes a 3x3 adjacency matrix.
  • The sampleCoordinate function is a helper function that samples a Coordinate by sampling the $x$, $y$, and $z$ values from normal distributions using normal.
  • The randomBonds function generates a random 3x3 adjacency matrix representing the bonds between atoms. Currently it is not implemented for brevity, but the implementation would be straightforward.
  • In the moleculeModel, after generating the molecule, we calculate the distance between the generated molecule and the observed molecule using the distMolecule function.
  • We use the score function to compute the likelihood of the distance given a normal distribution with mean 00 and standard deviation $1$. This scoring step conditions the model on the observed molecule, giving higher weights to generated molecules that are closer to the observed molecule.
  • The distMolecule function takes two molecules as input and calculates the distance between them. It extracts the coordinates of the atoms in each molecule and computes the Euclidean distance between all pairs of coordinates using euclideanDistance. The minimum distance among all pairs is returned as the distance between the molecules.
  • In the main function, we generate samples using Metropolis-Hastings and drop the first one thousand samples as a burn in. We then print the remaining molecules and weights.
By having an observed molecule and using the distMolecule function, we can condition our generative model on any molecule of interest. The scoring mechanism computes the distance between the generated molecule and the observed molecule, guiding the sampling process towards molecules that are more similar to the observed one.

### v3

@@ 3.2.1 Example of Orbitals : Hydrogen, Carbon & Oxygen
1-- Hydrogen atom: 1s^1
2hydrogen :: Shells
3hydrogen =
@@
6 , sSubShell = Just (SubShell
7 [ Orbital { orbitalType = So
8 , electronCount = 1
9 , orientation = Nothing
10 , hybridComponents = Nothing
11 }
12 ])
@@
18
19-- Carbon atom: 1s^2, 2s^2, 2p^2 (one electron in each of two p orbitals)
20carbon :: Shells
21carbon =
@@
24 , sSubShell = Just (SubShell
25 [ Orbital { orbitalType = So
26 , electronCount = 2
27 , orientation = Nothing
28 , hybridComponents = Nothing
29 }
30 ])
@@
37 , sSubShell = Just (SubShell
38 [ Orbital { orbitalType = So
39 , electronCount = 2
40 , orientation = Nothing
41 , hybridComponents = Nothing
42 }
43 ])
44 , pSubShell = Just (SubShell
45 [ Orbital { orbitalType = Px
46 , electronCount = 1
47 , orientation = Just (Coordinate 1 0 0)
48 , hybridComponents = Nothing
49 }
50 , Orbital { orbitalType = Py
51 , electronCount = 1
52 , orientation = Just (Coordinate 0 1 0)
53 , hybridComponents = Nothing
54 }
55 ])
@@
Listing 10: Examples of electronic configurations for hydrogen, carbon, and oxygen using the Orbital Algebraic Data Type representation in Haskell.
For the hydrogen atom, the electronic configuration is modeled using a single shell with a principal quantum number of 1. This shell contains only an $s$ subshell (represented by the sSubShell field) consisting of one orbital. The orbital is defined with the following properties:
  • orbitalType = So, indicating an $s$ orbital.
  • electronCount = 1, meaning it is occupied by one electron.
  • orientation = Nothing, which reflects the spherical symmetry of $s$ orbitals (they have no preferred spatial direction).
  • hybridComponents = Nothing, showing that the orbital is pure (i.e., it is not a hybrid orbital).
This precisely encodes the ground state configuration of hydrogen as 1s1.
For the carbon atom, the electronic configuration is represented by two shells. In the first shell (with principal quantum number 1), there is an $s$ subshell containing an orbital with:
  • orbitalType = So,
  • electronCount = 2 (representing 1s2),
  • orientation = Nothing, and
  • hybridComponents = Nothing.
In the second shell (with principal quantum number 2), the configuration comprises:
  • An $s$ subshell with an orbital defined similarly (with electronCount = 2 for the 2s electrons), and
  • A $p$ subshell (pSubShell) containing two orbitals:
    • The first $p$ orbital is defined as Orbital Px 1 (Just (Coordinate 1 0 0)), meaning it is of type Px, carries one electron, and has its orientation specified along the $x$-axis.
    • The second $p$ orbital is defined as Orbital Py 1 (Just (Coordinate 0 1 0)), meaning it is of type Py, carries one electron, and is oriented along the $y$-axis.
Both $s$ and $p$ orbitals in carbon have their hybridComponents fields set to Nothing, indicating that they are pure orbitals and not hybridized. This representation encodes the configuration 1s2 2s2 2p2, with one electron occupying each of the two available $p$ orbitals.
Overall, the inclusion of the orientation and hybridComponents fields in each orbital allows the data structure not only to capture the electron occupancy and energy level (via the principal quantum number) but also to describe the spatial characteristics of the orbitals. This detailed representation is fundamental for applications such as predicting molecular acidity, where orbital hybridisation and electron spatial distribution significantly influence chemical reactivity and intermolecular interactions.
3.3 Reactions
The proposed ADT can be extended to model chemical reactions. By defining a Reaction data type that captures the reactants, products, and conditions under which a reaction occurs, we can create a tool for studying and simulating chemical processes.
Listing 11 illustrates a Reaction data type, where reactants and products are represented as lists of pairs, containing a stoichiometric coefficient (a Double) and a Molecule. The conditions field captures reaction conditions such as temperature and pressure, defined with respect to a time range. Listing 12 provides an example reaction, illustrating the formation of water from hydrogen and oxygen under specific thermodynamic conditions.
Although these examples are simple, they demonstrate the flexibility of the ADT by extending it to model reaction processes. The compositional nature and modularity of the type system ensures that additional features not implemented here, such as catalysts or thermal energy, could be straightforwardly implemented. Stoichiometric or thermodynamic constraints could be enforced through use of dependent types, although this is left as future work.
1-- Reactions or transformations between chemical species
2data Reaction = Reaction
3 { reactants :: [(Double, Molecule)]
4 , products :: [(Double, Molecule)]
5 , conditions :: [Condition]
6 , rate :: Double
@@ 3.3 Reactions
8
9-- Conditions under which a reaction occurs
10data Condition = TempCondition {temperature :: Double}
11 | PressureCondition {presssure :: Double}
12
13data Times = Times { startTime :: Double, endTime :: Double }
Listing 11: Haskell representation of a chemical reaction including reactants, products, conditions, and reaction rate. The Reaction data type is composed using the Molecule data type
1-- Example reaction: 2H2 + O2 -> 2H2O
2exampleReaction :: Reaction
3exampleReaction = Reaction
4 { reactants = [(2.0, undefined), (1.0, undefined)]
5 , products = [(2.0, undefined)]
6 , conditions =
7 [ TempCondition 500.0
8 , PressureCondition 1.0
9 ]
10 , rate = 0.1
@@
Listing 12: Example reaction: formation of water from hydrogen and oxygen
In contrast, string-based representations such as reaction SMILES, SMARTS, and SMIRKS attempt to encode reactions as sequences of characters [12]. Reaction SMILES, for example, separates reactants and products using the “¿¿” operator, but does not inherently encode reaction conditions, catalysts, or rate dependencies [13]. SMARTS and SMIRKS provide pattern-matching capabilities for substructure searching and reaction templates, respectively, but lack a native mechanism for enforcing reaction validity [12].
String-based representations also inherit the limitations of the string data type itself. A lack of compositionality, and difficulties in formal reasoning make it unclear how one might extend SELFIES, for instance, to define a reaction SELFIES in which all syntactically valid strings correspond to chemically valid reactions [1]. Even if such an extension were possible, enforcing network-wide properties such as reversibility, equilibrium conditions, or pathway constraints would be nontrivial.
To generalise what is written here to adequately capture the semantics of reactions and reaction networks, it would be useful to derive an implementation off of a categorical semantics such as [63]. However, this is left as future work. Chemical reaction networks align naturally with category-theoretic formalisms, where reactions can be treated as morphisms between reactant and product sets [63]. This view has been used to model biochemical pathways, open dynamical systems, and Petri nets in systems chemistry [63], making an ADT in a typed functional language with support for categorical constructs such as Functors etc, an obvious choice for representation.
While constraints such as stoichiometric balance are not enforced in our implementation, the ADT provides a foundation upon which such constraints could be incorporated in future work through dependent types or additional verification layers.
However, the compositional nature of the ADT, its ability to be amended to incorporate reaction properties, and its alignment with category theoretic formalisms make it a natural framework cheminformatics applications.
@@ 3.4 Group Theory
In the following code snippet, a MolRot consists of a Molecule, a rotation axis specified by a Coordinate, and a rotation angle (in radians). The Group typeclass defines the necessary operations for a group, namely the group operation (mul), the inverse operation (inv), and the identity operation (e), and associativity can be proved but is implicit.
@@
The following provides an example of how a MoleculeRotation data type can be defined and made an instance of the Group typeclass, though the full implementational details are not included due to their complexity.
1class Group g where
2 mul :: g -> g -> g -- Group operation: combining two operations
3 inv :: g -> g -- Inverse of an operation
4 e :: g -> g -- Identity operation
5
6data MolRot = MolRot Molecule Coordinate Double deriving (Show)
7
8instance Group MolRot where
9 mul (MolRot mol1 axis1 angle1) (MoleculeRotation mol2 axis2 angle2) =
10 combineRotations mol1 axis1 angle1 axis2 angle2
11
12 inv (MoleculeRotation mol axis angle) = MolRot mol axis (-angle)
13
14 e (MoleculeRotation mol axis angle) = MolRot mol (Coordinate 0 0 0) 0
Listing 13: An instance of making a MolRot a group which has a rotation axis and a rotation angle.
@@ 3.5 Storage, Transmission, and Readability
As well the computational purposes discussed elsewhere in the article, the ADT provides advantages with respect to storage of the molecule, and transmission between cheminformaticians.
Molecules in the ADT can be serialised and stored as .hs files, which may be used for storage, and for transmission of molecular data, perhaps via large chemical databases, through straightforward use of the Show and Read instances in Haskell, as shown in Listing 15.
Additionally, as .hs files will be interpreted with Haskell, molecules can be straightforwardly compressed as Listing 14 demonstrates by using foldr to compress the bondsData list of the benzene molecule. Although this may raise the issue of “canonicalisation” for storage of molecules, it should be noted the bondsData in Listing 14 and Listing 8 will be stored the same way inside of a Haskell program, as the corresponding data is the same in each case. This could be beneficial in the representation of large molecules, as only a small number of bonds or atoms need be explicitly described to encode the molecule for storage.
1bondsData :: [((Int, Int), Bond)]
@@
Listing 14: Using foldr to compress the bondsData of the molecule benzene in the ADT representation. As Benzene is symmetric, the bonding system can be specified by only describing three bonding systems, and using those to construct the bondsData list.
@@
Listing 15: Haskell code demonstrating the serialisation and writing to disk of a Molecule data type and the reading of a molecule stored in an .hs file, through use of the show and read functions.
@@
3.6 A simple example of probablistic programming
Some code is excluded from listing 16 to conserve space, but is explained below.
This module implements a probabilistic generative model for molecules, where the goal is to generate a candidate molecule that is scored according to its similarity to an observed molecule. The observed molecule is obtained by reading an SDF file using the function observedMoleculeIO, which invokes parseDB1File to parse the file "./molecules/water.sdf". If no molecule is found, the function terminates with an error; otherwise, the first molecule is returned.
The core of the generative process is encapsulated in the function moleculeModel, which takes an observed Molecule as input and returns a computation in the Meas monad. Within this function, three atoms are generated by executing a monadic loop (via replicateM). For each atom, a chemical element is randomly selected from the set [C, N, O, H] using the uniformD distribution, and a three-dimensional coordinate is sampled from independent normal distributions by means of the helper function sampleCoordinate. The attributes associated with the chosen element, namely elementAttributes and elementShells, are then used to construct an Atom with a placeholder identifier (initially 0). After all atoms are generated, unique identifiers (1, 2, 3) are assigned to them.
Subsequently, the code computes all unique unordered pairs of atom identifiers (which, in the case of three atoms, yields the pairs (1,2), (1,3), and (2,3)) using list comprehensions and the function tails. For each pair, the code samples a Boolean value from a uniform distribution over [True, False] to decide whether a bond should be included between the corresponding atoms. When a bond is to be added, a bond order is sampled uniformly from the set [1, 2, 3] and is converted into a Bond record by setting its field delocNum to 2, 4, or 6, respectively. The bond is then stored in a map (of type M.Map (Integer, Integer) BondType) using a key that is the sorted pair of atom identifiers, thereby ensuring the symmetric nature of the bond representation.
Once both the atoms and bonds have been generated, a Molecule is constructed. Its quality is evaluated by computing the Hausdorff distance between the generated molecule and the observed molecule via the function hausdorffDistance. This distance is defined as the maximum over the minimum Euclidean distances between each atom of one molecule and the atoms of the other, where the Euclidean distance is computed by the function euclideanDistance. The resulting distance is then used to score the molecule by applying a normal probability density function (normalPdf) with mean 0 and standard deviation 1, through the score function.
The main function ties these components together. It first obtains the observed molecule using observedMoleculeIO, and then it executes a Metropolis–Hastings sampling procedure (via mh with a step size of 0.1) on the generative model instantiated with the observed molecule. After discarding an initial burn-in of 1000 samples, 1000 samples along with their corresponding weights are extracted and printed to standard output.
1moleculeModel :: Molecule -> Meas Molecule
2moleculeModel observed = do
@@ 3.6 A simple example of probablistic programming
4 atomsUnnumbered <- replicateM numAtoms \$ do
@@
8 shellsVar = elementShells symbol
9 return \$ Atom { atomID = 0, atomicAttr = attr, coordinate = coord, shells = shellsVar }
10
[…]
12 atomIDs = map atomID atoms
13
14 let possiblePairs = [(i, j) | (i:rest) <- tails atomIDs, j <- rest]
[…]
16 includeBond <- sample \$ uniformD [True, False]
17 if includeBond
18 then do
19 bondOrder <- sample \$ uniformD [1, 2, 3]
20 let bondType = case bondOrder of
21 1 -> Bond { delocNum = 2, atomIDs = Nothing }
22 2 -> Bond { delocNum = 4, atomIDs = Nothing }
23 3 -> Bond { delocNum = 6, atomIDs = Nothing }
24 _ -> error "Invalidbondorder"
25 return \$ Just (pair, bondType)
26 else return Nothing)
27 possiblePairs
28
29 let bondsList = catMaybes bondsMaybe
30 bondsMap = M.fromList bondsList
31 let molecule = Molecule { atoms = atoms, bonds = bondsMap }
32 let distance = hausdorffDistance molecule observed
33
@@
36
37sampleCoordinate :: Meas Coordinate
38sampleCoordinate = do
@@
43
44main :: IO ()
45main = do
46 observed <- observedMoleculeIO
47 samples <- mh 0.1 (moleculeModel observed)
48 let (molecules, weights) = unzip \$ take 1000 \$ drop 1000 samples
49 putStrLn \$ "Sampledmolecules:" ++ show molecules
50 putStrLn \$ "Weights:" ++ show weights
Listing 16: An example probabilistic program, where trace Metropolis-Hastings (M-H) operates by sampling entire “traces” of latent variables (i.e., all random choices) from a model [47]. Each step proposes a small perturbation to the current trace, controlled by a jitter parameter, often the standard deviation of the proposal distribution. This parameter must be tuned carefully to ensure efficient exploration of the posterior while maintaining a reasonable acceptance rate. Typically, an initial burn-in period is applied, during which early samples (which have not yet converged) are discarded. This burn-in helps remove the influence of the initial point on the chain, leading to samples that more accurately reflect the stationary distribution of the target posterior. Here the jitter is set to 0.1 and the number of burnin samples is 1000.
1euclideanDistance :: Coordinate -> Coordinate -> Double
2euclideanDistance (Coordinate x1 y1 z1) (Coordinate x2 y2 z2) =
@@
4
5hausdorffDistance :: Molecule -> Molecule -> Double
6hausdorffDistance mol1 mol2 =
@@
9 d1 = [minimum [euclideanDistance c1 c2 | c2 <- coords2] | c1 <- coords1]
10 d2 = [minimum [euclideanDistance c2 c1 | c1 <- coords1] | c2 <- coords2]
11 in max (maximum d1) (maximum d2)
Listing 17: Euclidean and Hausdorff distance calculations for molecular coordinates
In summary, the code defines a complete pipeline in which a molecule is generated by randomly selecting atom types, coordinates, and bonds, and the quality of the generated molecule is measured against an observed molecule using the Hausdorff distance. The sampling procedure then explores the space of possible molecules according to the probabilistic model defined.
3.6.1 Meas & Prob Monads
A monad is an abstraction that represents computations as sequences of steps, encapsulating context such as side effects, state, or randomness. In Haskell, a monad is defined by the operations return (to inject a value into the monadic context) and bind (represented as >>=), which chains computations while automatically managing the underlying context. This structure allows for composability and abstraction of computational effects, enabling programmers to write clean, modular code that implicitly threads additional information—be it state, failure, or randomness—throughout a computation.